### v3

# CAMEO: Collection of Multilingual Emotional Speech Corpora

Abstract
This paper presents CAMEO – a curated collection of multilingual emotional speech datasets designed to facilitate research in emotion recognition and other speech-related tasks. The main objectives were to ensure easy access to the data, to allow reproducibility of the results, and to provide a standardized benchmark for evaluating speech emotion recognition (SER) systems across different emotional states and languages. The paper describes the dataset selection criteria, the curation and normalization process, and provides performance results for several models. The collection, along with metadata, and a leaderboard, is publicly available via the Hugging Face platform.
Index Terms—  speech emotion recognition, multilingual dataset, benchmark
1 Introduction
In recent years, speech emotion recognition (SER) has gained attention due to its potential applications in human-computer interaction, mental health monitoring, and customer service. However, developing effective SER systems depends on high-quality datasets that capture the interplay of vocal cues and emotional states.
Despite the growing number of emotional speech corpora, the field still lacks standardization. Inconsistencies are evident in emotion taxonomies, transcription availability, and inclusion of speaker information. The metadata is often scattered across publications rather than linked to audio samples, which limits its usability. Moreover, the access to these datasets is also challenging. Unlike automatic speech recognition (ASR) or text-to-speech (TTS) corpora, which are commonly hosted on platforms like Hugging Face or GitHub, SER datasets are more difficult to locate, and some are shared without proper licensing.
These issues are exacerbated by the absence of standardized tools and benchmarks. Progress toward robust, multilingual SER systems is hindered by the fact that many studies evaluate models on a single dataset or language.
To address these challenges, we introduce CAMEO, a curated, publicly available collection of emotional speech datasets designed for reproducible benchmarking and multilingual evaluation.
The main contributions of this work include:
  • Release of the CAMEO collection on Hugging Face for transparent and reproducible SER benchmarking [^1] https://huggingface.co/datasets/amu-cai/CAMEO.
  • Systematic curation and normalization of 13 datasets spanning eight languages.
  • Comparative evaluation of multiple models on the unified corpus using standardized prompts and metrics.
  • A public leaderboard on Hugging Face for ongoing model comparison [^2] https://huggingface.co/spaces/amu-cai/cameo-leaderboard.
2 Related Work
One of the primary limitations of existing emotional speech datasets is their monolingual scope, which hinders the development of multilingual SER systems. Since most corpora cover only one language, the results obtained using them often fail to generalize across languages and cultures, where emotional expression varies.
Several studies have highlighted this issue. For example, Donuk et al. [7] reported 66.01% accuracy on the CREMA-D [2] dataset using convolutional neural networks with particle swarm optimization, while Catania et al. [4] achieved 82.34% accuracy on the Italian Emozionalmente dataset by fine-tuning wav2vec 2.0. Though promising, such results remain language-specific. Similarly, Mocanu et al. [17] showed that multimodal systems combining audio and video outperform audio-only approaches. On the RAVDESS dataset [14], for example, accuracy increased from 76% (audio only) to 83% (audio and video). On the CREMA-D dataset, accuracy increased from 62% to 82%. These results underscore the need for higher-quality audio resources to bridge the gap with multimodal systems.
The EmoBox [15] initiative attempts to address these challenges by aggregating multiple corpora into one repository, thereby easing the access but leaving the heterogeneity unresolved. Researchers must still download each dataset individually and navigate its unique characteristics, such as different sampling rates and metadata formats. Meanwhile, benchmarks such as EmoBench [21] evaluate emotional intelligence in LLMs using hand-crafted questions. However, these benchmarks remain text-only and exclude audio.
Our work builds on these efforts by focusing on standardizing emotional speech datasets and enabling the multilingual evaluation of SER systems through text instructions and audio prompts.
Fig. 1: Example illustrating the application of the post-processing strategy used to extract labels from generated responses.
3 Corpus Design
The CAMEO collection was designed to provide accessible, multilingual emotional speech data and ensure the reproducibility of evaluation results. The goal was to compile relevant emotion recognition corpora that could also be used for other speech-related research. To support cross-lingual and cross-cultural studies, we prioritized linguistic diversity by including a broad range of languages.
All corpora included in the collection are freely available under licenses that permit non-commercial use and derivative works. Licensing information is included in the metadata. To maximize usability, the collection is accompanied by extensive documentation and standardized metadata. Additionally, open-source evaluation code is provided to enable benchmarking and facilitate reproducibility.
Since the datasets are public and can be used for model training, the collection does not define train or test splits. This decision reflects the intended use of the collection as a benchmark for evaluating cross-lingual model performance. Moreover, taking into consideration the limited amount of data usable for SER, further splitting would reduce diversity and not solve the problem of contamination.
3.1 Selection Criteria for the Included Datasets
The inclusion of a dataset in the collection was determined by the following criteria:
  1. 1.
    Availability and License: The corpus is freely accessible and licensed for non-commercial use and derivative works.
  2. 2.
    Transcription: The dataset contains transcriptions, either directly or through associated publications, to increase its usability for other speech-related tasks.
  3. 3.
    Emotions: The annotations of basic emotional states are provided and consistent with standard taxonomies, such as Plutchik’s Wheel.
  4. 4.
    Metadata (optional): Information about speakers (IDs and demographics) was considered valuable. If such metadata was not available, relavant details were derived from documentation or publications.
3.2 Selected Corpora
The CAMEO collection consists of 13 carefully selected datasets, which are presented in Table 1.
Dataset Language No. samples
CaFE [11] French 936
CREMA-D [2] English 7 442
EMNS [18] English 1 205
Emozionalmente [3] Italian 6 902
eNTERFACE [16] English 1 257
JL-Corpus [12] English 2 400
MESD [9, 8] Spanish 862
nEMO [5] Polish 4 481
Oréau [13] French 502
PAVOQUE [22] German 5 442
RAVDESS [14] English 1 440
RESD [1] Russian 1 396
SUBESCO [23] Bengali 7 000
Table 1: Summary of the datasets included in the CAMEO collection.
The collection consists of a total of 41 265 audio samples. English accounts for approximately one-third of the material, reflecting its dominance in SER research. The collection represents 17 different emotions, with over 93% of samples annotated with one of the seven primary states: anger, disgust, fear, happiness, neutrality, sadness, and surprise.
Speaker-related metadata is widely available. Specifically, 94.5% of the samples include speaker identifiers, and 92.9% include gender annotations. Of these samples, 43.6% are spoken by female speakers, and 49.3% are spoken by male speakers.
3.3 Dataset Curation and Processing
The development of the CAMEO collection followed a structured, multi-stage process:
  1. 1.
    Data Acquisition: All datasets were downloaded from their original sources, inspected for consistency, assigned unique IDs, and purged of duplicates.
  2. 2.
    Standardization: Audio samples were converted to FLAC (16-bit, 16 kHz). Metadata and transcriptions were unified across datasets. Whisper Large v2 [19] was used to align transcriptions when necessary.
  3. 3.
    Metadata Serialization: Metadata was stored in JSON Lines (UTF-8) format and linked to sample IDs.
  4. 4.
    Distribution: The curated collection was uploaded to Hugging Face with proper credit given to the dataset authors. The metadata fields are summarized in Table 2.
Field Description
file_id Unique identifier of the audio sample.
audio File path, raw waveform, and sampling rate.
emotion Expressed emotional state.
transcription Orthographic transcription of the utterance.
speaker_id Unique speaker identifier.
gender Gender of the speaker.
age Age of the speaker.
dataset Dataset of origin.
language Primary language of the sample.
license Original dataset license.
Table 2: Overview of the metadata fields available in the CAMEO corpus, as published on the Hugging Face platform.
4 Evaluation
The evaluation protocol was designed based on the following principles:
  1. 1.
    To ensure reproducibility of the results, the evaluation includes easily accessible systems with audio modality.
  2. 2.
    To promote transparency and allow comparison across systems, evaluation results are publicly available on a leaderboard hosted on the Hugging Face platform.
  3. 3.
    To capture the performance of the models across different emotions and languages, the evaluation employs widely used metrics, such as macro-averaged F1 score, weighted F1 score and accuracy.
  4. 4.
    To enable further development, the benchmark is designed to be easily extensible. It allows for the easy integration of new datasets, metrics, methods, and systems.
To ensure consistency and fairness, all systems were evaluated using the same strategy. Each model was given a textual instruction and a single audio input. No additional metadata, such as speaker identity, gender, or language, was provided during inference.
The models were expected to produce a single-word output corresponding to the predicted emotional state. However, due to the generative nature of the evaluated systems, the models occasionally generated descriptive responses despite the given instructions. Additionally, the models tended to respond with an adjective rather than a noun.
As SER already poses a challenge, a post-processing strategy was designed to ensure that the models would not be penalized for minor errors. If a generated response is not an exact match for any of the labels, it is normalized and split into words. Then, the Levenshtein ratio between each target label and each word in the generated response is calculated. Similarity scores below a predefined threshold of 0.57 are filtered out for each label. The value of threshold was determined based on the Levenshtein ratio between the noun and adjective forms of the seven primary emotional states, as shown in Table 3. After filtering out the similarity scores, the remaining values are summed to yield an aggregated score for the given label. The label with the highest aggregated similarity score is selected as the best match. An example illustrating the application of the post-processing strategy is presented in Figure 1.
Noun Adjective Levenshtein Ratio
anger angry 0.8000
disgust disgusted 0.8750
fear fearful 0.7273
happiness happy 0.5714
neutrality neutral 0.8235
sadness sad 0.6000
surprise surprised 0.9412
Table 3: Levenshtein similarity ratios (LR) between the noun and adjective forms of the seven most frequently occuring emotions in the selected datasets.
Model Bengali English French German Italian Polish Russian Spanish All
Ichigo 0.7 0.1253 0.1293 0.1351 0.1226 0.1250 0.1489 0.1343 0.1634 0.1277
Qwen2 0.0 0.2508 0.3435 0.3154 0.3341 0.1913 0.1767 0.2450 0.2333 0.2023
SeaLLMs 0.3 0.0628 0.2168 0.1896 0.1216 0.0999 0.0547 0.0976 0.1761 0.1599
Ultravox 0.7 0.1163 0.1680 0.3073 0.2841 0.1383 0.1392 0.2086 0.2175 0.1484
Average 0.1388 0.2144 0.2369 0.2156 0.1386 0.1299 0.1714 0.1976 0.1596
Table 4: F1 macro scores obtained by each model, at a selected temperature setting, reported per language.
Model anger disgust fear happiness neutral sadness surprise
Ichigo 0.7 0.1414 0.1317 0.1366 0.1502 0.0377 0.2021 0.0793
Qwen2 0.0 0.1618 0.3211 0.1822 0.0367 0.0255 0.3905 0.4783
SeaLLMs 0.3 0.2246 0.2274 0.0153 0.0069 0.0135 0.2668 0.0956
Ultravox 0.7 0.2270 0.3420 0.0593 0.1673 0.0609 0.2193 0.1728
Average 0.1887 0.2556 0.0984 0.0903 0.0344 0.2697 0.2065
Table 5: F1 scores obtained for primary emotional states across the full corpus.
5 Benchmark Results
Four different models were evaluated: Qwen2-Audio-7B-Instruct [6], Ichigo-llama3.1-s-instruct-v0.4 [20], SeaLLMs-Audio-7B [24] and ultravox-v0_5-llama-3_1-8b [10]. To examine their behavior under different conditions, each system was evaluated using three different temperature settings: 0.0, 0.3, and 0.7.
Table 6 reports the overall performance using macro-averaged F1 score, weighted F1 score, and accuracy. Performance remained low across all models and temperatures, highlighting the difficulty of multilingual speech emotion recognition in a zero-shot setting. The best-performing system was Qwen2-Audio, which achieved a macro F1 score of 0.20 at a temperature of 0.0. Though modest, this result is consistent with the absence of task-specific fine-tuning and emphasizes the difficulty of mapping acoustic emotion cues to specific labels across languages using general-purpose AudioLLMs.
Model Macro F1 Weighted F1 Accuracy
Ichigo 0.0 0.1164 0.1171 0.1518
Ichigo 0.3 0.1164 0.1262 0.1549
Ichigo 0.7 0.1277 0.1359 0.1494
Qwen2 0.0 0.2023 0.3710 0.3911
Qwen2 0.3 0.1985 0.3704 0.3387
Qwen2 0.7 0.2005 0.3591 0.3825
SeaLLMs 0.0 0.1461 0.1887 0.2383
SeaLLMs 0.3 0.1599 0.1782 0.2200
SeaLLMs 0.7 0.1399 0.1616 0.2135
Ultravox 0.0 0.1417 0.2061 0.2333
Ultravox 0.3 0.1369 0.1972 0.2251
Ultravox 0.7 0.1484 0.2014 0.2294
Table 6: Overall evaluation results computed across all samples in the corpus.
Table 4 shows the macro F1 scores for each language at a fixed temperature. Qwen2-Audio consistently outperformed the other models across all evaluated languages, suggesting stronger generalization capabilities or broader exposure to emotionally annotated speech during pretraining. Qwen2-Audio demonstrated notably higher performance for English, French, and German. Ultravox also demonstrated higher performance for French and German, albeit to a lesser extent. These results suggest the possibility of data contamination, especially since these languages are well represented in publicly available emotional speech corpora.
To investigate potential overlap effects, macro F1 scores were computed at the dataset level. Qwen2-Audio performed substantially better on CREMA-D (0.80), RAVDESS (0.73), and eNTERFACE (0.54). A similar pattern was observed for Ultravox, which showed elevated performance on eNTERFACE (0.79) and Oréau (0.58). Since these datasets are widely used benchmarks in SER, the results suggest that they were likely encountered during training or fine-tuning.
Table 5 shows F1 scores for the sever primary emotional states. Sadness was the most reliably recognized emotion across all systems, while neutral speech was the most challenging. This may be due to the fact that sadness often exhibits more stable and distinctive acoustic cues, such as reduced pitch variation and slower tempo. In contrast, neutral speech is inherently heterogeneous and often acoustically overlaps with low-arousal emotional states. Despite its narrative or conversational nature, the lack of salient prosodic markers likely contributes to frequent misclassification.
6 Conclusions
This work introduces CAMEO, a publicly available collection of multilingual emotional speech datasets designed to facilitate transparent, reproducible benchmarking for SER. The corpus is released alongside evaluation tools and an open leaderboard for model comparison. The evaluation of several systems with the audio modality revealed low overall performance, highlighting the difficulty of the SER task, especially in multilingual and zero-shot settings.
Future work will focus on several key areas. First, the CAMEO collection will be expanded to include additional datasets, particularly those representing low-resource languages and underrepresented emotional states. Second, we intend to evaluate new models to make CAMEO a continuously developing resource for advancing SER research.